OpenStudy (anonymous):

I swear if I do not get the answer soon I will have a fluttering panic attack :) Which ordered pairs in the form (x, y) are solutions to the equation 3x - 4y = 21? Choose all answers that are correct. A. 3, 3 B. 1,6 C. 7, 0 D. 11, 3

OpenStudy (texaschic101):

this is just a matter of subbing and seeing which ones are true. 3x - 4y = 21 A.) (3,3)...x = 3 and y = 3 now sub 3(3) - 4(3) = 21 9 - 12 = 21 -3 = 21 (incorrect...so this is not a solution) B.) (1,6)...x = 1 and y = 6 now sub again 3(1) - 4(6) = 21 3 - 24 = 21 -21 = 21 (incorrect...not this one either) C.) (7,0)...x = 7 and y = 0 3(7) - 4(0) = 21 21 - 0 = 21 21 = 21 (This is correct...this is a solution) D.) (11,3)...x = 11 and y = 3 3(11) - 4(3) = 21 33 - 12 = 21 21 = 21 (correct...this is also a solution) So the solutions to the equations are : C and D